A personal computer manufacturer buys 43 % of its chips from Japan and the rest from America. 2 % of the Japanese chips are defective and 1.1 % of the American chips are defective.

1. Find the probability that a chip is defect-free.
2. Find the probability that a chip is defective and made in Japa

Answer by robertb(5830)

P(defect-free) = P(defect-free AND from Japan) + P(defect-free AND from America) = = P(defect-free|from Japan)*P(from Japan) + P(defect-free| from America)*P(from America) = 0.98*0.43 + 0.989*0.57 = 0.98513. Or 98.513%.
P(defective AND from Japan) = P(defect|from Japan)*P(from Japan) = 0.02*0.43 = 0.0086, or 0.86 %.